1.Given X and n, find p and q. Show your calculation steps.





1.X=56; n=80



2.X=35; n=96



3.X=420; n=1000



4.X=636; n=1200



5.X=1200 n=10000

1
Expert's answer
2022-02-02T05:06:32-0500

1.


"p=\dfrac{x}{n}=\dfrac{56}{80}=\dfrac{7}{10}=0.7"

"q=1-p=1-0.7=0.3"


2.


"p=\dfrac{x}{n}=\dfrac{35}{96}"

"q=1-p=1-\dfrac{35}{96}=\dfrac{61}{96}"

3.


"p=\dfrac{x}{n}=\dfrac{420}{1000}=\dfrac{21}{50}=0.42"

"q=1-p=1-0.42=0.58"



4.


"p=\dfrac{x}{n}=\dfrac{636}{1200}=\dfrac{53}{100}=0.53"

"q=1-p=1-0.53=0.47"

5.


"p=\dfrac{x}{n}=\dfrac{1200}{10000}=\dfrac{3}{25}=0.12"

"q=1-p=1-0.12=0.88"